### v2

@@ I Introduction
The $T_{cc}^{+}$ tetraquark falls into the family of doubly heavy tetraquarks, formed out of two heavy quarks and two light antiquarks. Doubly heavy tetraquarks were predicted to be stable under QCD interactions in the heavy-quark mass limit [8, 9, 10]. A crucial question is whether the bottom quark is sufficiently heavy for the $bb\bar{u}\bar{d}$ tetraquark to form a bound state below the $BB^{*}$ two-meson threshold. Several phenomenological investigations utilizing Quark level Models [11, 8, 12, 13, 14, 15, 16, 17, 18, 19, 10, 20, 21, 22, 23, 24, 25, 26, 27], QCD sum rules [28, 29, 30], effective field theory [31, 32, 33, 34] and Chromomagnetic Interaction Model [35, 36, 37, 38, 39] suggest that this is indeed the case. There is a long way to go for the experimental detection of doubly bottom hadrons considering the center of momentum energy it would require to produce a simultaneous pair of bottom quark-antiquark combination. Strategies for the experimental detection of tetraquarks with bottomness-$2$ are discussed in Refs. [40, 41, 42].
Numerical investigations based on first principles, such as those using Lattice Quantum Chromodynamics (lattice QCD) provide an excellent tool in studying the properties of these exotic states. Early lattice calculations explored the $bb\bar{u}\bar{d}$ tetraquark treating the bottom quarks in the static limit under the Born-Oppenheimer approximation within a quenched approximation [43] and with only light dynamical quark flavors [44, 45]. Recently, dynamical simulations involving light and strange quark dynamics have become more common with b-quark evolution more often approximated using a nonrelativistic QCD (NRQCD) evolution on the lattice [46, 47, 48, 49, 50, 51, 52, 53]. Our earlier work also employed the NRQCD framework [47]. Collectively, these studies provide compelling evidence that the isoscalar axialvector $bb\bar{u}\bar{d}$ channel should host a strong interaction stable bound state tetraquark. In the past few years, lattice investigations have also extended to other heavy tetraquark flavor combinations involving bottom quarks. For instance, systems such as $bc\bar{u}\bar{d}$ have been analyzed in our previous work [54, 55], and the $bb\bar{u}\bar{s}$ has been explored in [48, 50]. While there is qualitative agreement on the existence of stable axialvector doubly bottom tetraquarks in the isovector $bb\bar{u}\bar{d}$ and in the isodoublet $bb\bar{u}\bar{s}$ channels, the inferences arrived from different lattice investigations differ in the $bc\bar{u}\bar{d}$ case.
@@
The study of isoscalar $bs\bar{u}\bar{d}$ channels is a continuation of our previous calculations of a similar tetraquark system with flavor content $bc\bar{u}\bar{d}$ [54, 55]. Only a single lattice QCD study exists addressing this channel [56]. Several previous studies using QCD sum rules and model calculations have indicated the possibility of bound states for $bs\bar{u}\bar{d}$ with $I=0$ in these channels. For instance, the quark color delocalization screening model predicts bound states in both quantum channels [57]. Similarly, the SU(3) chiral quark model [58] and an alternative chiral quark model [22] support the existence of bound states, although the latter suggests a weaker binding. A QCD sum rule analysis of the scalar bottom-strange channel [59] points to a deeply bound state. However, an early study using a non-chiral model [60] found no evidence for a bound state in the $I(J^{P})=0(1^{+})$ channel. The only existing lattice study indicates a weakly bound state for this system and did not include a finite volume analysis [56]. Naturally, further lattice investigations are required to reach definitive conclusions in these channels.
@@
In the valence sector, we utilize the same setup that was used in our previous investigations [63, 47, 64, 54, 55]. In this framework, an overlap fermion formulation [65, 66] is used to describe the dynamics of valence quark flavors with masses up to the charm quark mass. We compute valence quark propagators for four different quark masses in the quark mass range with representative pseudoscalar meson masses $M_{ps}\sim$ 0.5, 0.6, 0.7, and 3.0 GeV. A landscape plot of the pseudoscalar meson masses studied across the four lattice QCD ensembles can be found in Figure 1 of Ref. [55]. The quark propagators with $M_{ps}\sim$ 0.7 and 3.0 GeV correspond to the physical strange and the charm quarks. The valence strange quark mass was tuned to reproduce the mass of hypothetical $\eta_{s}$ meson ($m_{\eta_{s}}=688.5$ MeV) [67], which cannot mix with the physical $\eta$ and $\eta^{\prime}$ mesons via the self annihilation of $s\bar{s}$. The valence charm quark mass was tuned by setting the kinetic mass of the charmonium 1S spin average $\{a\overline{M}_{kin}^{\bar{c}c}=0.75aM_{kin}(J/\psi)+0.25aM_{kin}(\eta_{c})\}$ on each ensemble studied to the physical charmonium 1S spin average mass (3068 MeV) [68]. Similar to most other lattice studies of heavy quarkonium, we neglect the OZI suppressed heavy quark self annihilation diagrams in the tuning procedure.
@@ III Measurements
Here $\{\Phi_{i}(\mathbf{x},t)\}$ is a carefully crafted basis of interpolating currents that has the set of quantum numbers of the state we are interested in. The information on the quantum numbers of the state is encoded in the operator-state overlap $Z_{i}^{n}=\bra{0}\Phi_{i}\ket{n}$, the strength of which determines the coupling of the interpolator, $\Phi_{i}$, with the state $n$. The mass or energy of the state can be extracted from the large time behavior of these correlation functions.
@@ VII Scattering amplitudes
Following the extraction of ground state energies, we extract the $S$-wave amplitudes in the channels studied within elastic approximation following the finite-volume two-particle spectrum quantization prescription á la Lüscher and its generalizations [89, 90]. For all the three cases we investigate, within an elastic scattering assumption the $S$-wave phase shift $\delta_{l=0}(k)$ can be extracted from finite-volume spectral energies via the quantization relation $p\cot[\delta_{0}(p)]=2Z_{00}[1;(\frac{pL}{2\pi})^{2})]/(L\sqrt{\pi})$, where $p$ is the momentum of the scattering particles in their center of momentum frame. $p$ is related to the total energy $E_{cm}$ through $4sp^{2}=(s-(M_{1}+M_{2})^{2})(s-(M_{1}-M_{2})^{2})$, where $s=E_{cm}^{2}$ is the Mandelstam variable and $M_{i}$ is the mass of the scattering particle. We enforce the quantization constraint to extract the energy dependence of the amplitude following the procedure discussed in Appendix B of Ref. [81]. Following the extraction of the amplitude, we search for possible near threshold poles representing any bound states in the amplitude. Given the constraints on going beyond a constant parametrization in $p^{2}$, we would not be able to access any information on resonances that might exist. A real bound state exists in the $S$-wave amplitude, when $p{\mathrm{cot}}\delta_{0}=+\sqrt{-p^{2}}$.
@@ VII.1 $bb\bar{u}\bar{d}$ tetraquarks
where $A^{[0]}=-1/a_{0}$ is the negative of inverse scattering length in the continuum limit. The lattice spacing dependence of the extracted $p~{}\cot(\delta_{0})$ in units of $E_{BB^{*}}$ is presented in Figure 8, whereas the continuum estimate is presented along with the respective lattice degrees of freedom in the standard $p~{}\cot(\delta_{0})$ versus $p^{2}$ format in Figure 9. The crossing of the continuum extrapolated amplitude with the (real/virtual) bound state pole constraint curves given by the (orange/cyan) parabolic curve indicates the position of any below threshold pole. These subthreshold pole positions for each $M_{ps}$ are represented by the magenta hexagons. In Figure 10, we present the continuum extrapolated amplitudes as a function of the $M_{ps}$ involved. Note that the amplitude is stable with respect to the $M_{ps}$ across the lowest three values, indicating hardly any dependence on the light meson mass involved in the $BB^{*}$ interactions. The invariance of the amplitude on the $M_{ps}$ involved suggests negligible effects on the chiral extrapolated amplitude with different fit forms, such as a constant or a quadratic in $M_{ps}$ inspired from chiral effects or a linear form inspired from Heavy Quark Effective theory, compared to the respective statistical uncertainties. The resultant chiral extrapolated amplitude suggests a bound state pole in the physical $BB^{*}$ amplitude with a binding energy of $\Delta E_{T_{bb}}(1^{+})=-116(^{+30}_{-36})$ MeV. One may argue the largeness of $M_{ps}$ values considered in such chiral extrapolations. The variation in this estimate if one had chosen the amplitude at the lightest $M_{ps}$ we investigate or with a naïve continuum extrapolation of the binding energy estimates at the lowest $M_{ps}$, is well within a few MeV about 116 MeV with reasonable reduced chi-square values between 1 and 1.5. In the future, we hope to extend our investigations to lighter $M_{ps}$ values to further quantify this. We also observe that the amplitude and the resultant binding energy estimate at the physical point are hardly influenced by the inclusion of the finite-volume data at $M_{ps}\sim 3.0$ GeV.
@@ VII.2 $bs\bar{u}\bar{d}$ tetraquarks
In Figures 13 and 14, we present the results for $BK$ scattering based on similar amplitude fits as discussed in the previous cases. Similar to the axialvector $bs\bar{u}\bar{d}$ tetraquarks, the continuum extrapolated amplitudes are consistently zero indicating infinite scattering length. The large uncertainties in the amplitudes also lead to large errors in the binding energies extracted from those. All of these are transparent enough to be readily inferred from Figures 13 and 14. In short, the $bs\bar{u}\bar{d}$ finite-volume data at hand does not support any statistically significant signatures for a bound axialvector or scalar $bs\bar{u}\bar{d}$ tetraquark.
@@ VIII Summary and Conclusions
In Tables 6 and 5 and Figure 15, we present a compilation of various estimates for the binding energy of $T_{bb}$ tetraquarks following lattice as well as nonlattice methodologies, together with the estimate we arrive from this investigation. No lattice determination, including this work, unambiguously suggests the existence of $bs\bar{u}\bar{d}$ bound state, whereas a few phenomenological approaches argue possible binding. More dedicated methods, incorporating large statistics and a rigorous finite-volume analysis, are essential to reaching a definitive conclusion for this system, given its proximity to the threshold. We intend to undertake such a study in the future. We collect existing predictions for this system in Figure 16 and Table 7. We remark that there is only one other lattice investigation of the $bs\bar{u}\bar{d}$ system [56], where the author claims no evidence for any bound $bs\bar{u}\bar{d}$ state. Since there is no estimate for the ground state energy provided in Ref. [56] for the $bs\bar{u}\bar{d}$ system, and hence is missing in Figure 16.
@@
In summary, our lattice setup suggests a deeply bound tetraquark in the axialvector $bb\bar{u}\bar{d}$ system with binding energy $\mathcal{O}(100$ MeV)))), whereas an unbound $bs\bar{u}\bar{d}$ system in both axialvector and scalar quantum channels. Note that a relatively shallow yet deeply bound state with axial-vector and scalar quantum numbers for the for $bc\bar{u}\bar{d}$ system, with a binding energy of approximately 40 MeV, was recently reported using the same lattice setup employed here [54, 55]. This result helps bridge the sequence of these tetraquarks, which have valence quark contents ranging from bottom to strange. While our inferences on the $bb\bar{u}\bar{d}$ system and the $bs\bar{u}\bar{d}$ system turn out to be consistent with other existing lattice determinations c.f. Ref. [48], the disagreements in binding energies for the $bc\bar{u}\bar{d}$ system with other existing lattice determination [82] remains to be understood, particularly considering the effect of discretization in these heavy hadrons.
Ref. [Year)] $N_{f}$ $a~{}[fm]$ $m^{sea}_{ps}$ [MeV] $S_{q}^{sea}$ $S_{c}^{val}$ Amplitude $\Delta E_{T_{bb}}(1^{+})$
@@
[43] (2012) 2 0.079 340 Twisted mass Wilson-TM Potential −$-57(19)$
[45] (2015) 2 0.042,0.079 340-352 Wilson-TM Wilson-TM Potential -90(40)
[46] (2017) 2+1 0.0899 164-415 Wilson - No -189(10)
[47] (2019) 2+1+1 0.0582-0.1207 153-689 HISQ Overlap No -143(34)
[48] (2019) 2+1 0.0828-0.1141 139-431 Domain wall Domain-wall LĂźscher-based -128(34)
[49] (2020) 2+1 0.09-0.15 - Asqtad HISQ No -189(18)
[50] (2023) 2+1 0.085 220-420 Wilson-Clover Wilson-Clover No -112(13)
[51] (2023) 2+1 0.064, 0.0907 416-701 Wilson-Clover Wilson-Clover HALQCD -155(17)(20)
@@
[52] (2024) 2+1+1 0.0872-0.1510 217-313 HISQ Wilson-clover No −$-100(10)(^{+36}_{-43})$
[53] (2024) 2+1 0.08 186-794 Wilson-Clover Wilson-clover No −115(17)11517-115(17)- 115 ( 17 )
This work 2+1+1 0.0582-0.1207 217-319 HISQ Overlap Lüscher-based −$-116(^{+30}_{-36})$
Table 5: Binding energy $\Delta E_{T_{bb}}(1^{+})$ predictions from various lattice QCD studies. The results are presented along with some of the relevant technical details pertaining to the respective lattice investigations. Two numbers associated with Ref. [51] correspond to the estimates from an elastic analysis (top) and a coupled $BB^{*}$-$B^{*}B^{*}$ analysis (bottom).
@@
[58] (2018) SU(3) chiral quark Model -70.2 -64.5

### v3

@@ I Introduction
The $T_{cc}^{+}$ tetraquark falls into the family of doubly heavy tetraquarks, formed out of two heavy quarks and two light antiquarks. Doubly heavy tetraquarks were predicted to be stable under QCD interactions in the heavy-quark mass limit [8, 9, 10]. A crucial question is whether the bottom quark is sufficiently heavy for the $bb\bar{u}\bar{d}$ tetraquark to form a bound state below the $BB^{*}$ two-meson threshold. Several phenomenological investigations utilizing Quark level Models [11, 8, 12, 13, 14, 15, 16, 17, 18, 19, 10, 20, 21, 22, 23, 24, 25, 26, 27, 28], QCD sum rules [29, 30, 31], effective field theory [32, 33, 34, 35] and Chromomagnetic Interaction Model [36, 37, 38, 39, 40] suggest that this is indeed the case. There is a long way to go for the experimental detection of doubly bottom hadrons considering the center of momentum energy it would require to produce a simultaneous pair of bottom quark-antiquark combination. Strategies for the experimental detection of tetraquarks with bottomness-$2$ are discussed in Refs. [41, 42, 43].
Numerical investigations based on first principles, such as those using Lattice Quantum Chromodynamics (lattice QCD) provide an excellent tool in studying the properties of these exotic states. Early lattice calculations explored the $bb\bar{u}\bar{d}$ tetraquark treating the bottom quarks in the static limit under the Born-Oppenheimer approximation with only light dynamical quark flavors [44, 45]. Recently, dynamical simulations involving light and strange (and charm in some studies) quark dynamics have become more common with b-quark evolution more often approximated using a nonrelativistic QCD (NRQCD) evolution on the lattice [46, 47, 48, 49, 50, 51, 52, 53]. Our earlier work also employed the NRQCD framework [47]. Collectively, these studies provide compelling evidence that the isoscalar axialvector $bb\bar{u}\bar{d}$ channel should host a strong interaction stable bound state tetraquark. In the past few years, lattice investigations have also extended to other heavy tetraquark flavor combinations involving bottom quarks. For instance, systems such as $bc\bar{u}\bar{d}$ have been analyzed in our previous work [54, 55], and the $bb\bar{u}\bar{s}$ has been explored in [48, 50]. While there is qualitative agreement on the existence of stable axialvector doubly bottom tetraquarks in the isovector $bb\bar{u}\bar{d}$ and in the isodoublet $bb\bar{u}\bar{s}$ channels, the inferences arrived from different lattice investigations differ in the $bc\bar{u}\bar{d}$ case. A summary and more detailed account of various lattice investigations of doubly heavy tetraquarks can be found in the Refs. [56, 57].
@@
The study of isoscalar $bs\bar{u}\bar{d}$ channels is a continuation of our previous calculations of a similar tetraquark system with flavor content $bc\bar{u}\bar{d}$ [54, 55]. Only a single lattice QCD study exists addressing this channel [58]. Several previous studies using QCD sum rules and model calculations have indicated the possibility of bound states for $bs\bar{u}\bar{d}$ with $I=0$ in these channels. For instance, the quark color delocalization screening model predicts bound states in both quantum channels [59]. Similarly, a few recent quark model investigations [60, 22, 61] support the existence of bound states, although the latter suggests a weaker binding. A QCD sum rule analysis of the scalar bottom-strange channel [62] points to a deeply bound state. However, an early study using a non-chiral model [63] found no evidence for a bound state in the $I(J^{P})=0(1^{+})$ channel. The only existing lattice study indicates a weakly bound state for this system and did not include a finite volume analysis [58]. Naturally, further lattice investigations are required to reach definitive conclusions in these channels.
@@
In the valence sector, we utilize the same setup that was used in our previous investigations [66, 47, 67, 54, 55]. In this framework, an overlap fermion formulation [68, 69] is used to describe the dynamics of valence quark flavors with masses up to the charm quark mass. We compute valence quark propagators for four different quark masses in the quark mass range with representative pseudoscalar meson masses $M_{ps}\sim$ 0.5, 0.6, 0.7, and 3.0 GeV. A landscape plot of the pseudoscalar meson masses studied across the four lattice QCD ensembles can be found in Figure 1 of Ref. [55]. Note that the chosen $M_{ps}$ values are rather heavier than the physical point and it is more desirable to have further closer-to-physical $M_{ps}$ values. However, the available statistics at lighter $M_{ps}$ values turns out to be insufficient to extract results with reasonable statistical precision. The quark propagators with $M_{ps}\sim$ 0.7 and 3.0 GeV correspond to the physical strange and the charm quarks. The valence strange quark mass was tuned to reproduce the mass of hypothetical $\eta_{s}$ meson ($m_{\eta_{s}}=688.5$ MeV) [70], which cannot mix with the physical $\eta$ and $\eta^{\prime}$ mesons via the self annihilation of $s\bar{s}$. The valence charm quark mass was tuned by setting the kinetic mass of the charmonium 1S spin average $\{a\overline{M}_{kin}^{\bar{c}c}=0.75aM_{kin}(J/\psi)+0.25aM_{kin}(\eta_{c})\}$ on each ensemble studied to the physical charmonium 1S spin average mass (3068 MeV) [71]. Similar to most other lattice studies of heavy quarkonium, we neglect the OZI suppressed heavy quark self annihilation diagrams in the tuning procedure.
@@ III Measurements
Here $\{\Phi_{i}(\mathbf{x},t)\}$ is a carefully crafted basis of interpolating currents that has the set of quantum numbers of the state we are interested in. The mass or energy of the state can be extracted from the large time behavior of these correlation functions. The strength of operator-state overlap $Z_{i}^{n}=\bra{0}\Phi_{i}\ket{n}$ determines the coupling of the interpolator $\Phi_{i}$ with the state $n$.
@@ VII Scattering amplitudes
Following the extraction of ground state energies, we extract the $S$-wave amplitudes in the channels studied within elastic approximation following the finite-volume two-particle spectrum quantization prescription á la Lüscher and its generalizations [93, 94]. For all the three cases we investigate, within an elastic scattering assumption the $S$-wave phase shift $\delta_{l=0}(k)$ can be extracted from finite-volume spectral energies via the quantization relation $p\cot[\delta_{0}(p)]=2Z_{00}[1;(\frac{pL}{2\pi})^{2}]/(L\sqrt{\pi})$, where $p$ is the momentum of the scattering particles in their center of momentum frame. $p$ is related to the total energy $E_{cm}$ through $4sp^{2}=(s-(M_{1}+M_{2})^{2})(s-(M_{1}-M_{2})^{2})$, where $s=E_{cm}^{2}$ is the Mandelstam variable and $M_{i}$ is the mass of the scattering particle. We enforce the quantization constraint to extract the energy dependence of the amplitude following the procedure discussed in Appendix B of Ref. [85]. Following the extraction of the amplitude, we search for possible near threshold poles representing any bound states in the amplitude. Given the constraints on going beyond a constant parametrization in $p^{2}$, we would not be able to access any information on resonances that might exist. A real bound state exists in the $S$-wave amplitude, when $p{\mathrm{cot}}\delta_{0}=+\sqrt{-p^{2}}$.
@@ VII.1 $bb\bar{u}\bar{d}$ tetraquarks
where $A^{[0]}=-1/a_{0}$ is the negative of inverse scattering length in the continuum limit. The lattice spacing dependence of the extracted $p~{}\cot(\delta_{0})$ in units of $E_{BB^{*}}$ is presented in Figure 8, whereas the continuum estimate is presented along with the respective lattice degrees of freedom in the standard $p~{}\cot(\delta_{0})$ versus $p^{2}$ format in Figure 9. The crossing of the continuum extrapolated amplitude with the (real/virtual) bound state pole constraint curves given by the (orange/cyan) parabolic curve indicates the position of any below threshold pole. These subthreshold pole positions for each $M_{ps}$ are represented by the magenta hexagons. In Figure 10, we present the continuum extrapolated amplitudes as a function of the $M_{ps}$ involved. Note that the amplitude is stable with respect to the $M_{ps}$ across the lowest three values, indicating hardly any dependence on the light meson mass involved in the $BB^{*}$ interactions. The invariance of the amplitude on the $M_{ps}$ involved suggests negligible effects on the chiral extrapolated amplitude with different fit forms, such as a constant or a quadratic in $M_{ps}$ inspired from chiral effects or a linear form inspired from Heavy Quark Effective theory, compared to the respective statistical uncertainties. The resultant chiral extrapolated amplitude suggests a bound state pole in the physical $BB^{*}$ amplitude with a binding energy of $\Delta E_{T_{bb}}(1^{+})=-116(^{+30}_{-36})$ MeV. One may argue the largeness of $M_{ps}$ values considered in such chiral extrapolations. The variation in this estimate if one had chosen the amplitude at the lightest $M_{ps}$ we investigate or with a naïve continuum extrapolation of the binding energy estimates at the lowest $M_{ps}$, is well within a few MeV about 116 MeV with reasonable reduced chi-square values between 1 and 1.5. Note that the observed independence of the scattering length with $M_{ps}$ might be a consequence of the inability to resolve the valence quark mass dependence with the available statistics, and towards the chiral limit there could be systematic uncertainities that one might be blind about. In the future, we hope to extend our investigations to lighter $M_{ps}$ values to further quantify this. We also observe that the amplitude and the resultant binding energy estimate at the physical point are hardly influenced by the inclusion of the finite-volume data at $M_{ps}\sim 3.0$ GeV.
@@ VII.2 $bs\bar{u}\bar{d}$ tetraquarks
In Figures 13 and 14, we present the results for $BK$ scattering based on similar amplitude fits as discussed in the previous cases. Similar to the axialvector $bs\bar{u}\bar{d}$ tetraquarks, the continuum extrapolated amplitudes are consistently zero indicating infinite scattering length. The large uncertainties in the amplitudes also lead to large errors in the binding energies extracted from those. All of these are transparent enough to be readily inferred from Figures 13 and 14. In short, the $bs\bar{u}\bar{d}$ finite-volume data at hand does not support any statistically significant signatures for a bound axialvector or scalar $bs\bar{u}\bar{d}$ tetraquark. Similar to the arguments in stated in the case of $bb\bar{u}\bar{d}$, future lattice simulations closer to the physical pion mass $M_{\pi}$ are essential to extend our findings towards the chiral limit.
@@ VIII Summary and Conclusions
In Tables 6 and 5 and Figure 15, we present a compilation of various estimates for the binding energy of $T_{bb}$ tetraquarks following lattice as well as nonlattice methodologies, together with the estimate we arrive from this investigation. We observe our new estimate, that involved a conservative determination of the ground state energies followed by a finite-volume scattering analysis, is consistent with all the predictions from other lattice QCD determinations, except two outliers with deeper binding. Owing to the presence of heavy quarks, the finite-volume binding energy estimates in this heavy tetraquark system could be affected by discretization effects. Ours is the only calculation that has utilized finite-volume data at three lattice spacings that go as fine as ∼$\sim 0.06$ fm, and also attempt a finite-volume analysis in the $T_{bb}$ system. Our work utilizes a chiral quark action for the light, strange and charm quark fields in contrast to most other existing studies, which commonly uses a Wilson-clover action, offers a complementary perspective. Moreover, using the same fermion action across a range of systems provides a valuable consistency check on the binding energies in these tetraquark systems.
Despite the relatively heavy pseudoscalar meson masses $M_{ps}$ used and the lack of meson-meson operators with separately momentum-projected mesons in this study, our estimate seem to agree very well with existing lattice determinations that include them. Although this is encouraging, future investigations with $M_{ps}$ values closer to the physical point, incorporating local as well as bilocal meson-meson interpolators, accompanied by a rigorous finite-volume analysis, are essential to reach at rigorous and conservative estimates on the bindings also accommodating for various unaccounted systematics in the systems addressed here.
@@
No lattice determination, including this work, unambiguously suggests the existence of $bs\bar{u}\bar{d}$ bound state, whereas a few phenomenological approaches argue possible binding. We collect existing predictions for this system in Figure 16 and Table 7. We remark that there is only one other lattice investigation of the $bs\bar{u}\bar{d}$ system [58], where the author claims no evidence for any bound $bs\bar{u}\bar{d}$ state. Since there is no estimate for the ground state energy provided in Ref. [58] for the $bs\bar{u}\bar{d}$ system, and hence is missing in Figure 16. For similar reasons, as stated in the above paragraph, future lattice simulations closer to the physical point and with interpolating operators that mimic bilocal meson-meson structures are crucial in assessing the interactions towards the chiral limit and to reaffirm or to rule out our findings.
@@
In summary, our lattice setup suggests a deeply bound tetraquark in the axialvector $bb\bar{u}\bar{d}$ system with binding energy $\mathcal{O}(100$ MeV)))), whereas an unbound $bs\bar{u}\bar{d}$ system in both axialvector and scalar quantum channels. Note that a relatively shallow yet deeply bound state with axial-vector and scalar quantum numbers for the $bc\bar{u}\bar{d}$ system, with a binding energy of approximately 40 MeV, was recently reported using the same lattice setup employed here [54, 55]. This result helps bridge the sequence of these tetraquarks, which have valence quark contents ranging from bottom to strange. While our inferences on the $bb\bar{u}\bar{d}$ system and the $bs\bar{u}\bar{d}$ system turn out to be consistent with other existing lattice determinations c.f. Ref. [48], the disagreements in binding energies for the $bc\bar{u}\bar{d}$ system with other existing lattice determination [86] remains to be understood, particularly considering the effect of discretization in these heavy hadrons.
Ref. [Year)] $N_{f}$ $a~{}[fm]$ $n_{a}$ $m^{sea}_{ps}$ [MeV] $S_{q}^{sea}$ $S_{q}^{val}$ $S_{b}^{val}$ Amplitude $\Delta E_{T_{bb}}(1^{+})$
@@
[44] (2012) 2 0.079 1 340 Twisted mass Twisted mass Static Potential −$-57(19)$
[45] (2015) 2 0.042,0.079 2 340, 352 Twisted mass Twisted mass Static Potential -90(40)
[46] (2017) 2+1 0.0899 1 164-415 Wilson-clover Wilson-clover NRQCD No -189(10)
[47] (2019) 2+1+1 0.0582-0.1207 3 153-689 HISQ Overlap NRQCD No -143(34)
[48] (2019) 2+1 0.0828-0.1141 3 139-431 Domain wall Domain-wall NRQCD LĂźscher-based -128(34)
[49] (2020) 2+1 0.09-0.15 3 - Asqtad HISQ NRQCD No -189(18)
[50] (2023) 2+1 0.085 1 220-420 Wilson-Clover Wilson-Clover NRQCD No -112(13)
[51] (2023) 2+1 0.0907 1 416-701 Wilson-Clover Wilson-Clover NRQCD HALQCD -155(17)(20)
@@
[52] (2024) 2+1+1 0.0872-0.1510 3 217-313 HISQ Wilson-clover NRQCD No −$-100(10)(^{+36}_{-43})$
[53] (2024) 2+1 0.08 1 164-707 Wilson-Clover Wilson-clover NRQCD No −115(17)11517-115(17)- 115 ( 17 )
This work 2+1+1 0.0582-0.1207 3 217-319 HISQ Overlap NRQCD Lüscher-based −$-116(^{+30}_{-36})$
Table 5: Binding energy $\Delta E_{T_{bb}}(1^{+})$ predictions from various lattice QCD studies. The results are presented along with some of the relevant technical details pertaining to the respective lattice investigations. $n_{a}$ refers to the number of distinct lattice spacings used in each calculation and the bold faced numbers are cases, where a continuum extrapolation was performed. Two numbers associated with Ref. [51] correspond to the estimates from an elastic analysis (top) and a coupled $BB^{*}$-$B^{*}B^{*}$ analysis (bottom).
@@
[60] (2018) SU(3) chiral quark Model -70.2 -68